# Homeomorphisms of continua through projective Fraïssé limits

### v1

@@ 1. Introduction
As an application of our methods based on projective Fraïssé theory, we prove the following theorem that is phrased purely in terms of the homeomorphism group of the pseudoarc. (We provide some information on the pseudoarc in the introduction. A stronger version of the theorem below is proved as Theorem 4.1.)
@@
The statement above was conjectured for $n=1$ by Kwiatkowska [11] and was established, again in the case $n=1$, by Bice and Malicki [2]. We prove the general statement for arbitrary $n$, in fact, for $\mathbb{N}$}, using an argument based on projective Fraïssé theory. Our methods are different from those of [2]. They also appear to be simpler.
@@
where $I=n\geq 1$ or $I={\mathbb{N}}$. This line of investigation was initiated by Glasner and Weiss [6]; groups whose diagonal conjugacy actions with $n=1$ had a dense orbit were said to possess the topological Rohlin property in that paper. A general theory behind the existence of dense orbits for diagonal conjugacy actions for automorphism groups of countable structures was developed by Kechris and Rosendal in [8]. A survey of the area can be found in [5]. The reader may consult the papers [1], [2], [4], [5], [6], [8], [10], [11], [12], [15] for examples of Polish groups with dense orbits in diagonal conjugacy actions.
@@
However, the goal of this paper is broader than merely proving Theorem 1.1—we investigate the relationship between the automorphism groups of projective Fraïssé limits and the homeomorphism groups of compact spaces that are canonical quotients of those limits. The mathematical context for this investigation is as follows. One starts with a countable (up to isomorphism) family $\mathcal{K}$} of finite reflexive graphs and a family of epimorphisms among graphs in $\mathcal{K}$}. When writing $\mathcal{K}$} we have in mind the family of reflexive graphs together with the family of epimorphisms. The binary reflexive graph relation on the structures in $\mathcal{K}$} is denoted by $R$. Assuming that the epimorphisms in $\mathcal{K}$} fulfill a projective amalgamation condition, a canonical projective limit $\mathbb{K}$ of $\mathcal{K}$} exists. The object $\mathbb{K}$ is a compact totally disconnected metric space that can be equipped with a canonical interpretation $R^{\mathbb{K}}$ of $R$. This is a compact binary reflexive graph relation on $\mathbb{K}$, and with a set of continuous epimorphisms from $\mathbb{K}$} to the structures in $\mathcal{K}$. If $R^{\mathbb{K}}$ is transitive, then it is a compact equivalence relation, which allows us to form the quotient space $K={\mathbb{K}}/R^{\mathbb{K}}$, which is a compact metric space. Even though $\mathbb{K}$ as a topological space is totally disconnected, the quotient space $K$ is often connected, that is, it is a continuum; in fact, this is the most interesting situation from the topologically point of view.
@@
We consider two groups associated with the objects above, the automorphism group ${\rm Aut}({\mathbb{K}})$ of $\mathbb{K}$} and the homeomorphism group ${\rm Homeo}(K)$ of the quotient space $K$. Both these groups come with natural Polish group topologies on them. Furthermore, there is a natural continuous homomorphism
${\rm pr}\colon{\rm Aut}({\mathbb{K}})\to{\rm Homeo}(K).$
Elements of ${\rm Aut}({\mathbb{K}})$ are easier to deal with than elements of ${\rm Homeo}(K)$ as they are essentially combinatorial objects. Likewise, the group ${\rm Aut}({\mathbb{K}})$ is easier to deal with than the group ${\rm Homeo}(K)$, for example, it is non-archimedean. We exploit this drop in complexity when passing from ${\rm Homeo}(K)$ to ${\rm Aut}({\mathbb{K}})$ in order to study ${\rm Homeo}(K)$ through ${\rm Aut}({\mathbb{K}})$. For more precise information on projective Fraïssé theory, see Appendix A.
To achieve this goal, we need to consider the topology on ${\rm Aut}({\mathbb{K}})$ that is obtained by pulling back the topology on ${\rm Homeo}(K)$ using the continuous homomorphism $\rm pr$}. By continuity of $\rm pr$, this topology is weaker than the natural topology on ${\rm Aut}({\mathbb{K}})$. In Section 2, we give a combinatorial description of the topology on ${\rm Aut}({\mathbb{K}})$ inherited from ${\rm Homeo}(K)$ through $\rm pr$}.
@@
Then we move to finding combinatorial conditions corresponding to the existence of dense orbits under diagonal conjugacy actions. Since we will be assuming that ${\rm Aut}({\mathbb{K}})$ is dense in ${\rm Homeo}(K)$, when considering density of diagonal conjugacy actions, it suffices to consider conjugacy by elements of ${\rm Aut}({\mathbb{K}})$. This leads to three types of conjugacy actions. Given $n\in{\mathbb{N}}$, $n\geq 1$, we consider:
@@
    conjugacy by elements of ${\rm Aut}({\mathbb{K}})$ of tuples in ${\rm Aut}({\mathbb{K}})^{n}$;
@@
    conjugacy by elements of ${\rm Aut}({\mathbb{K}})$ of tuples in ${\rm Aut}({\mathbb{K}})^{n}$ taken with the topology inherited from ${\rm Homeo}(K)^{n}$;
@@
    there exist comeagerly many $\overline{\gamma}\in{\rm Aut}({\mathbb{K}})^{n}$ with conjugacy orbits dense in ${\rm Aut}({\mathbb{K}})^{n}$;
@@
    there exist comeagerly many $\overline{\gamma}\in{\rm Aut}({\mathbb{K}})^{n}$ with conjugacy orbits dense in ${\rm Homeo}(K)^{n}$;
@@
The three combinatorial conditions are related but distinct. They are versions of the Joint Projection Property for suitable categories and are obtained by dualizing to the projective setting and adapting to the mix of two topologies (coming from ${\rm Aut}({\mathbb{K}})$ and ${\rm Homeo}(K)$) of the Joint Embedding Property in the paper by Kechris and Rosendal [8]. Our conditions are stated, and the theorems establishing the equivalence between the appropriate Joint Projection Property and the existence of dense orbits in the corresponding conjugacy action are proved in Section 3. We note that case (b)/(b) exhibits the most interesting interaction between the two topologies on ${\rm Aut}({\mathbb{K}})$: comeagerness refers to the natural topology on ${\rm Aut}({\mathbb{K}})$, while density of orbits concerns the topology inherited from ${\rm Homeo}(K)$. This is precisely the case applied to the pseudoarc. Finally, note that since the property of a point having a dense orbit under a continuous action of a Polish group is $G_{\delta}$, the properties in (a) and (c) are equivalent to the existence of a single dense orbit in ${\rm Aut}({\mathbb{K}})^{n}$ and ${\rm Homeo}(K)^{n}$, respectively.
For our applications, recall that the pseudoarc is a continuum, that is, a compact connected metric space, constructed by Knaster [9] and characterized by Bing [3] as the unique chainable hereditarily indecomposable continuum. In the same paper [3], Bing gave another characterization of the pseudoarc. It is the unique continuum that is generic in the following sense. Let $\mathcal{C}$} be the space of all continua included in the Hilbert cube $0,1]^{\mathbb{N}}$ endowed with the Vietoris topology, which is induced by the Hausdorff metric. The space $\mathcal{C}$} is a compact metric space. It turns out [3] that there exists a continuum $P$ such that the subset of $\mathcal{C}$} consisting of copies of $P$ is comeager in $\mathcal{C}$}, in fact, it is a dense $G_{\delta}$. This continuum is the pseudoarc. For more information on the pseudoarc, the reader may consult the survey [13] or the book [14].
@@
where $n\in{\mathbb{N}}$. Morphisms in $\mathcal{P}$} are all epimorphisms among structures in $\mathcal{P}$}. As proved in [7], the family $\mathcal{P}$} forms a transitive projective Fraïssé class and its limit $\mathbb{P}$} is such that $P={\mathbb{P}}/R^{{\mathbb{P}}}$ is homeomorphic to the pseudoarc.
In Section 4, we apply the results of Section 3 to the pseudoarc. We prove the appropriate Joint Projection Property for the class $\mathcal{P}$}. This allows us to apply the results of Section 3 to show that there there exists an element of ${\rm Aut}({\mathbb{P}})^{\mathbb{N}}$ whose orbit with respect to the diagonal conjugacy action ${\rm Aut}({\mathbb{P}})$ on ${\rm Aut}({\mathbb{P}})^{\mathbb{N}}$ is dense in ${\rm Homeo}(P)^{\mathbb{N}}$; in fact, the set of such tuples is comeager in ${\rm Aut}({\mathbb{P}})^{\mathbb{N}}$. In particular, Theorem 1.1 follows.
Finally, in Section 5, we turn to the following natural problem concerning the conjugacy action of ${\rm Homeo}(P)$ on itself: is every homeomorphism in ${\rm Homeo}(P)$ conjugate in ${\rm Homeo}(P)$ to an element of ${\rm Aut}({\mathbb{P}})$? We answer it in the negative by exhibiting a homeomorphism whose conjugacy class misses ${\rm Aut}({\mathbb{P}})$. This is the first example of this phenomenon in the projective Fraïssé context. The construction uses our analysis of topologies in Section 2. The homeomorphism is constructed by finding a sequence of automorphisms in ${\rm Aut}({\mathbb{P}})$ that is Cauchy with respect to the uniformity inducing the topology on ${\rm Aut}({\mathbb{P}})$ that is inherited from ${\rm Homeo}(P)$ through the map $\rm pr$.
@@ 1. $\mathcal{K}$ is a countable projective Fraïssé family;
  2. $\mathbb{K}$ with the binary relation $R^{\mathbb{K}}$ is the projective Fraïsse limit of $\mathcal{K}$;
  3. if $R^{\mathbb{K}}$ is transitive, then $K=\mathbb{K}/R^{\mathbb{K}}$ is the quotient topological space and ${\rm pr}:\mathbb{K}\to K$ is the canonical projection.
@@
We will often identify ${\rm Aut}({\mathbb{K}})$ with its image under the homomorphism $\rm pr$} in ${\rm Homeo}(K)$, that is, we will consider $f\in{\rm Aut}({\mathbb{K}})$ as an element of ${\rm Homeo}(K)$ keeping in mind the identification spelled out in (A.1) in Appendix A. Note that this constitutes some abuse of terminology since, in general, $\rm pr$ need not be injective.

### v2

@@ 1. Introduction
As an application of our methods based on projective Fraïssé theory, we prove the following theorem that is phrased purely in terms of the homeomorphism group of the pseudoarc. (We provide some information on the pseudoarc later in the introduction.)
@@
Let $P$ be the pseudoarc, and let ${\rm Homeo}(P)$ be the group of all homeomorphisms of $P$ equipped with the uniform convergence topology.
@@
The statement above was conjectured for $n=1$ by Kwiatkowska [13] and was established, again for $n=1$, by Bice and Malicki [2]. We prove the general statement for arbitrary $n$, and also for $\mathbb{N}$}, using an argument based on projective Fraïssé theory. Our methods differ from those of [2] and also appear to be simpler. Theorem 1.1 follows from the more refined Theorem 1.2. In fact, in light of Theorem 1.3, Theorem 1.2 yields a stronger conclusion even in the case $n=1$, compared with [2]—see the discussion following Theorem 1.2.
@@
where $I=n\geq 1$ or $I={\mathbb{N}}$. This line of investigation was initiated by Glasner and Weiss [7]. When $G={\rm Homeo}(X)$ is the group of all homeomorphisms of a compact metric space $X$, the existence of dense orbits of the diagonal conjugacy actions translates into a statement about continuous actions of free groups on $X$ as follows. Let $\Gamma$ be a countable group. A continuous action of $\Gamma$ on $X$ is identified with a homomorphism from $\Gamma$ to $G$, so it is an element of $G^{\Gamma}$. Consequently, the space of all continuous actions of $\Gamma$ on $X$ is a subset of $G^{\Gamma}$. The homeomorphism group $G$ acts naturally by conjugation on this space of continuous action of $\Gamma$ on $X$. This type of framework was investigated, for example, in [4] and [5]. In this context, Theorem 1.1 amounts to asserting that the space of continuous actions of the free group $\mathbb{F}_{I}$ on the pseudoarc has a dense orbit, where $I=n\in{\mathbb{N}}$, $n>0$, or $I={\mathbb{N}}$. A general theory behind the existence of dense orbits for diagonal conjugacy actions for automorphism groups of countable structures was developed by Kechris and Rosendal in [10]. A survey of the area can be found in [6]. The reader may consult the papers [1], [2], [5], [6], [7], [10], [12], [13], [14], [17] for examples of Polish groups with dense orbits in diagonal conjugacy actions.
@@
The goal of this paper is broader than merely proving Theorem 1.1—we investigate the relationship between the automorphism groups of projective Fraïssé limits and the homeomorphism groups of compact spaces that are canonical quotients of those limits. The mathematical context for this investigation is as follows. (For detailed information on projective Fraïssé theory, see Appendix A.) One starts with a countable (up to isomorphism) family $\mathcal{K}$} of finite reflexive graphs and a family of epimorphisms among graphs in $\mathcal{K}$}. (A reflexive graph is a binary symmetric reflexive relation. An epimorphism is a function from the vertices of one reflexive graph to the vertices of another that preserves the edge relation and is surjective on edges and so, by reflexivity, also on vertices.) When writing $\mathcal{K}$} we have in mind the family of reflexive graphs together with the family of epimorphisms. The reflexive graph relation on the structures in $\mathcal{K}$} is denoted by $R$. Assuming that the epimorphisms in $\mathcal{K}$} fulfill a projective amalgamation condition, a canonical projective limit $\mathbb{K}$ of $\mathcal{K}$} exists. The object $\mathbb{K}$ is a compact totally disconnected metric space equipped with a set of continuous functions from $\mathbb{K}$} to the reflexive graphs in $\mathcal{K}$, with the reflexive graphs in $\mathcal{K}$} carrying the discrete topology. This structure induces a canonical interpretation $R^{\mathbb{K}}$ of $R$ on $\mathbb{K}$, which is a compact reflexive graph relation. If $R^{\mathbb{K}}$ is transitive, then it is a compact equivalence relation. In that situation, we form the quotient space $K={\mathbb{K}}/R^{\mathbb{K}}$, which is a compact metric space. Even though $\mathbb{K}$ as a topological space is totally disconnected, the quotient space $K$ is often connected, that is, it is a continuum—in fact, this is the most interesting situation from the topologically point of view.
@@
We consider two groups associated with the objects above, the automorphism group ${\rm Aut}({\mathbb{K}})$ of $\mathbb{K}$} and the homeomorphism group ${\rm Homeo}(K)$ of the quotient space $K$. Both these groups come with natural Polish group topologies on them, namely, the topologies of uniform convergence. Furthermore, there is a natural continuous homomorphism
(1.1) ${\rm pr}\colon{\rm Aut}({\mathbb{K}})\to{\rm Homeo}(K).$
We sometimes identify elements of ${\rm Aut}({\mathbb{K}})$ with their images under $\rm pr$ in ${\rm Homeo}(K)$. For more information on this identification, see the section Notation and Conventions below in the introduction.
Elements of ${\rm Aut}({\mathbb{K}})$ are easier to deal with than elements of ${\rm Homeo}(K)$, as they are essentially combinatorial objects and their properties are closely related to combinatorial properties of the family $\mathcal{K}$}. Thus, there is a drop in complexity when passing from ${\rm Homeo}(K)$ to ${\rm Aut}({\mathbb{K}})$. (Theorem 1.3 below is an example of a statement substantiating this claim.) We exploit this reduction in complexity in order to study ${\rm Homeo}(K)$ through ${\rm Aut}({\mathbb{K}})$, and ultimately through $\mathcal{K}$}. This is implemented in three steps. First, we transfer the uniform convergence topology on ${\rm Homeo}(K)$ to ${\rm Aut}({\mathbb{K}})$ using the continuous homomorphism $\rm pr$ in (1.1) and we give a combinatorial description of this transferred topology. Second, dualizing [10], we develop a combinatorial approach to the existence of dense orbits in the projective setting of ${\rm Aut}({\mathbb{K}})$ and ${\rm Homeo}(K)$. A new aspect of this development is the consideration of two distinct topologies on ${\rm Aut}({\mathbb{K}})$, which leads to three distinct versions of orbit density. And third, we give applications to the homeomorphism group of the pseudoarc and to the automorphism group of the pre-pseudoarc.
Thus, our first step is to consider the topology on ${\rm Aut}({\mathbb{K}})$ that is obtained by pulling back the topology on ${\rm Homeo}(K)$ using the continuous homomorphism $\rm pr$} from (1.1). Since we will be using both these topologies on the group ${\rm Aut}({\mathbb{K}})$, we introduce notation for each of them. Let
$\tau_{u}\;\hbox{ and }\;\tau_{i}$
stand, respectively, for the uniform convergence topology and the topology inherited from ${\rm Homeo}(K)$ via the homomorphism $\rm pr$ in (1.1) on ${\rm Aut}({\mathbb{K}})$. We also refer to the product topologies on ${\rm Aut}({\mathbb{K}})^{n}$, with $n\in{\mathbb{N}}$, $n\geq 1$, as $\tau_{u}$ and $\tau_{i}$. By continuity of $\rm pr$, the topology $\tau_{i}$ is weaker than $\tau_{u}$. In Section 2, Theorem 2.3 and Corollaries 2.4 and 2.5, we give combinatorial descriptions of the two topologies $\tau_{u}$ and $\tau_{i}$ on ${\rm Aut}({\mathbb{K}})$.
@@
After that, we move to finding combinatorial conditions corresponding to the existence of dense orbits under diagonal conjugacy actions. We note that the group ${\rm Homeo}(K)$ will always be considered with its natural topology (corresponding to uniform convergence on $K$) and the products ${\rm Homeo}(K)^{n}$ and ${\rm Homeo}(K)^{\mathbb{N}}$ with the products of the uniform convergence topology. Since we will be assuming that ${\rm Aut}({\mathbb{K}})$ is dense in ${\rm Homeo}(K)$, when considering density of orbits of diagonal conjugacy actions, it suffices to consider conjugacy by elements of ${\rm Aut}({\mathbb{K}})$. This leads to three types of conjugacy actions. Given $n\in{\mathbb{N}}$, $n\geq 1$, we consider:
@@
    conjugacy by elements of ${\rm Aut}({\mathbb{K}})$ of tuples in ${\rm Aut}({\mathbb{K}})^{n}$ taken with $\tau_{u}$;
@@
    conjugacy by elements of ${\rm Aut}({\mathbb{K}})$ of tuples in ${\rm Aut}({\mathbb{K}})^{n}$ taken with $\tau_{i}$;
@@
    there exist $\tau_{u}$-comeagerly many $\overline{\gamma}\in{\rm Aut}({\mathbb{K}})^{n}$ with conjugacy orbits $\tau_{u}$-dense in ${\rm Aut}({\mathbb{K}})^{n}$;
@@
    there exist $\tau_{u}$-comeagerly many $\overline{\gamma}\in{\rm Aut}({\mathbb{K}})^{n}$ with conjugacy orbits $\tau_{i}$-dense in ${\rm Aut}({\mathbb{K}})^{n}$;
@@
The three combinatorial conditions are related but distinct. They are versions of the Joint Projection Property for suitable categories and are obtained by dualizing to the projective setting and adapting to the mix of two topologies (one coming from ${\rm Aut}({\mathbb{K}})$ and the other one from ${\rm Homeo}(K)$) of the Joint Embedding Property in the paper by Kechris and Rosendal [10]. Our conditions are stated, and the theorems establishing the equivalence between the appropriate Joint Projection Property and the existence of dense orbits in the corresponding conjugacy action are proved in Section 3, Theorem 3.4 and Corollaries 3.10 and 3.11. We note that the case (b)/(b) exhibits the most interesting interaction between the two topologies on ${\rm Aut}({\mathbb{K}})$: comeagerness in (b) refers to the uniform convergence topology $\tau_{u}$ on ${\rm Aut}({\mathbb{K}})$, while density of orbits concerns the topology $\tau_{i}$ inherited from ${\rm Homeo}(K)$. This is precisely the case applied to the pseudoarc—see below. Finally, note that since the property of a point having a dense orbit under a continuous action of a Polish group is $G_{\delta}$, the properties in (a) and (c) are equivalent to the existence of a single dense orbit in ${\rm Aut}({\mathbb{K}})^{n}$ and ${\rm Homeo}(K)^{n}$, respectively.
For our applications, recall that the pseudoarc is a continuum, that is, a compact connected metric space, constructed by Knaster [11] and characterized by Bing [3] as the unique chainable hereditarily indecomposable continuum. In the same paper [3], Bing gave another compelling characterization of the pseudoarc. It is the unique continuum that is generic in the following sense. Let $\mathcal{C}$} be the space of all continua included in the Hilbert cube $0,1]^{\mathbb{N}}$ endowed with the Vietoris topology, that is, the topology induced by the Hausdorff metric. The space $\mathcal{C}$} is a compact metric space. As proved in [3], there exists a continuum $P$ such that “topologically most” elements of $\mathcal{C}$} are homeomorphic to $P$, that is, the subset of $\mathcal{C}$} consisting of copies of $P$ is comeager in $\mathcal{C}$}, in fact, it is a dense $G_{\delta}$. This continuum $P$ is the pseudoarc. For more information on the pseudoarc, the reader may consult the survey [15] or the book [16].
@@
where $n\in{\mathbb{N}}$. Morphisms in $\mathcal{P}$} are all epimorphisms among structures in $\mathcal{P}$}. As proved in [9], the family $\mathcal{P}$} forms a transitive projective Fraïssé class and its limit $\mathbb{P}$} is such that $P={\mathbb{P}}/R^{{\mathbb{P}}}$ is homeomorphic to the pseudoarc. In Section 4, we use this representation of the pseudoarc from [9] to apply our results of Section 3 to the pseudoarc. We prove the appropriate Joint Projection Property for the class $\mathcal{P}$}, which by the results of Section 3, yields the theorem below. A still stronger version of this result is stated and proved as Theorem 4.1.
Theorem 1.2.
For each natural number $n\geq 1$, there exists an element of ${\rm Aut}({\mathbb{P}})^{n}$ whose orbit with respect to the diagonal conjugacy action of ${\rm Aut}({\mathbb{P}})$ is dense in ${\rm Homeo}(P)^{n}$. Furthermore, there exists an element of ${\rm Aut}({\mathbb{P}})^{\mathbb{N}}$ whose orbit with respect to the diagonal conjugacy action ${\rm Aut}({\mathbb{P}})$ is dense in ${\rm Homeo}(P)^{\mathbb{N}}$.
Theorem 1.1 follows immediately from Theorem 1.2. Moreover, we note that Theorem 1.2 provides additional information compared to Theorem 1.1, as it asserts that the element with a dense conjugacy class lies in ${\rm Aut}({\mathbb{P}})^{n}$ or ${\rm Aut}({\mathbb{P}})^{\mathbb{N}}$, rather than merely in ${\rm Homeo}(P)^{n}$ or ${\rm Homeo}(P)^{\mathbb{N}}$. For $n=1$, this yields an additional improvement over the theorem proved in [2]. Of course, this is an improvement only if one can show that not every homeomorphism of $P$ can be conjugated to an element of ${\rm Aut}({\mathbb{P}})$. We now turn to this issue.
Since ${\rm Aut}({\mathbb{P}})$ is dense in ${\rm Homeo}(P)$, each homeomorphism of $P$ can be approximated by an automorphism of $\mathbb{P}$} to an arbitrary degree of precision. A natural problem is whether a homeomorphism of $P$ can in fact be realized as an automorphism of $\mathbb{P}$}. This problem can be stated precisely in terms of the conjugacy action of ${\rm Homeo}(P)$ on itself: is every homeomorphism in ${\rm Homeo}(P)$ conjugate, within ${\rm Homeo}(P)$, to an element of ${\rm Aut}({\mathbb{P}})$? This question is further motivated by the remarks following Theorem 1.2 above. In Section 5, we answer it in the negative by exhibiting a homeomorphism whose conjugacy class does not intersect ${\rm Aut}({\mathbb{P}})$. The following result is restated and proved as Theorem 5.1.
Theorem 1.3.
There exists an element of ${\rm Homeo}(P)$ that is not conjugate in ${\rm Homeo}(P)$ to an element of ${\rm Aut}({\mathbb{P}})$.
Theorem 1.3 above is the first example of the phenomenon described in it in the projective Fraïssé context. The construction uses our analysis of topologies in Section 2. The homeomorphism of $P$ in the conclusion of this theorem is obtained as the limit of a sequence of automorphisms in ${\rm Aut}({\mathbb{P}})$ that is Cauchy with respect to the uniformity inducing the topology $\tau_{i}$ on ${\rm Aut}({\mathbb{P}})$ inherited from ${\rm Homeo}(P)$. We expect this method to be useful in constructing homeomorphisms of $P$ with other properties.
@@ 1. $\mathcal{K}$ is a countable projective Fraïssé family;
  2. $\mathbb{K}$ with the binary relation $R^{\mathbb{K}}$ is the projective Fraïssé limit of $\mathcal{K}$;
  3. if $R^{\mathbb{K}}$ is transitive, then $K=\mathbb{K}/R^{\mathbb{K}}$ is the canonical quotient topological space and ${\rm pr}:\mathbb{K}\to K$ is the canonical projection.
@@
We often identify elements of ${\rm Aut}({\mathbb{K}})$ with their images under the homomorphism $\rm pr$} in ${\rm Homeo}(K)$, that is, we consider $f\in{\rm Aut}({\mathbb{K}})$ as the element ${\rm pr}(f)$ of ${\rm Homeo}(K)$ keeping in mind the identification spelled out in (A.2) in Appendix A. Such a move is justified by $\rm pr$ being injective under mild assumptions on $\mathcal{K}$}, see Proposition A.5. In particular, this is true in all the topological situations considered in the literature, for example, in the case of the pseudoarc and the associated with it projective Fraïssé family $\mathcal{P}$}.